METAL-MOLECULE INTERACTIONS IN In^{+}-NH_{2}(CH_{3}), In^{+}-NH(CH_{3})_{2}, In^{+}-N(CH_{3})_{3} and THEIR NEU-TRAL RADICALS

GRETCHEN K. ROTHSCHOPF, SHENGGANG LI AND \underline{DONG-SHENG YANG}, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The indium-amine radicals were produced by reactions of laser-vaporized metal atoms with methylamines seeded in helium gas and identified with time-of-flight mass spectrometry. Metal-ligand interactions of the positive ion and neutral species were probed using ZEKE spectroscopy. From the ZEKE spectra, ionization energies and metal-ligand stretching frequencies were determined for the three complexes, and bending frequencies were obtained for the monomethylamine and dimethylamine species. In comparison with InNH_{3},^{a} these methylated complexes have lower ionization energies and stronger metal-ligand binding.

aG. K. Rothschopf, J. S. Perkins, S. Li and D. -S. Yang, J. Phys. Chem. A \underline{104}, 8178 (2000).